OpenStudy (anonymous):

Find the rate of change with equation 5x + 3y = -2

OpenStudy (anonymous):

ok, I;m just confused where to start.

OpenStudy (amoodarya):

first find y =mx+b then m= delta y / delta x

OpenStudy (anonymous):

So am I solving for Y?

OpenStudy (amoodarya):

\[5x + 3y = -2 \\3y =-5x-2 \\\\y=\frac{ 1 }{3}(-5x-2)\\y=\frac{ -5x }{3}+\frac{ -2 }{3}\\ \frac{ \Delta y }{ \Delta x } =\frac{ f(x+ \Delta x)-f(x) }{ x+\Delta x -x }=\frac{ -5 }{ 3 }\]

OpenStudy (anonymous):

Ok, I did the first and second parts like you did. Just wondering why the 3rd part changes to a +. Is it because of the function?

OpenStudy (amoodarya):

yes line is a function like y=f(x)

OpenStudy (anonymous):

Ok, thank so much!